Your lab TA asks you to prepare a buffer solution with a pH of 7.40. The following reagents are all available to you: 550 mL of 0.200 M formic acid, 450 mL of 0.200 M hypochlorous acid, solid sodium formate and solid sodium hypochlorite. a) Which 2 ingredients will you use to prepare the desired buffer? b) What is the base/acid ratio in the desired buffer? c) What mass (in g) of solid base must you dissolve in the corresponding acid solution to achieve the desired buffer pH? d) If 1.00 mL of 1.00 M NaOH is added to the buffer solution, what will be the new pH?
Answer #1

Answer:

To make a buffer of desired pH, the value of pKa should be near to the pH of buffer solution to maximize the buffer capacity.

Here, the pKa of hypochlorous (HOCl) acid is 7.53.

a) We are going to use the HOCl and sodium hypochlorite solution.

b) Hence, for the preparation of buffer with pH = 7.40

pH = pKa + log(base/acid)

7.40 = 7.53 + log(base/acid) => log(base/acid) = -0.13

base/acid = 10-0.13 => base/acid = 0.7413

Base = 0.7413 acid

c) If acid is 0.2 M then, base = 0.14826 M

Molar mass of NaOCl = 74.44 g/mol and Total volume of solution = 450 mL

So, 0.14826 mol = 1000 mL

450 mL of solution contains = 0.066717 moles

Amount of NaOCl = 74.44 x 0.066717 = 4.9664 g of NaOCl is needed to add in 450 mL of 0.2 M of HOCl solution.

d) 1 mL of 1.0 M of NaOH contains 10-3 moles of OH- ions, which on addition produces the 10-3 moles of OCl- ions by neutralizing the similar amount of HOCl

Now moles of HOCl and NaOCl present = 0.089 moles and 0.067717 moles

Total volume = 451 mL or 0.451 L

and the concentrations are [HOCl] = 0.1973 M and [OCl-] = 0.1501 M

pH = 7.53 + log(0.1501/0.1973) = 7.41

Please let me know, if you have any doubts by commenting below the answer.

Thanks